# SpectrumWorld: Artificial Intelligence Foundation for Spectroscopy

### v3

@@ Abstract
Deep learning holds immense promise for spectroscopy, yet research and evaluation in this emerging field often lack standardized formulations. To address this issue, we introduce SpectrumLab, a pioneering unified platform designed to systematize and accelerate deep learning research in spectroscopy. SpectrumLab integrates three core components: a comprehensive Python library featuring essential data processing and evaluation tools, along with leaderboards; an innovative SpectrumAnnotator module that generates high-quality benchmarks from limited seed data; and SpectrumBench, a multi-layered benchmark suite covering 14 spectroscopic tasks and over 10 spectrum types, featuring spectra curated from over 1.2 million distinct chemical substances. Thorough empirical studies on SpectrumBench with 18 cutting-edge multimodal LLMs reveal critical limitations of current approaches. We hope SpectrumLab will serve as a crucial foundation for future advancements in deep learning-driven spectroscopy. The anonymous code and experimental records are available at https://ai4s-chem.github.io/SpectrumWorld/
@@ 1 Introduction
Spectroscopy, which investigates the interaction between electromagnetic radiation and matter, provides a powerful way to investigate the molecular structure and properties (2004; 2025). By capturing characteristic patterns, such as peaks and shifts, in signals analogous to audio waveforms, spectroscopy offers a compact, information-rich representation of molecular systems (2020). This low-dimensional encoding is indispensable in chemistry (2016; 2017; 2020), materials science (2025; 2025), and life sciences (2020; 2023; 2025). It is not only central to molecular structure elucidation (i.e., Spectrum-to-Molecule structure) and property prediction, but also a key enabler for new material discovery and drug screening. In recent years, machine learning methods, especially deep learning, have demonstrated tremendous potential in spectroscopic data analysis, opening a new era of automation and intelligence in spectroscopy research (2017a; 2019; 2020; 2022; 2023; 2024; 2025).
@@
Despite recent advances, deep learning for spectroscopy still faces several fundamental challenges. Specifically, high-quality experimental spectral data remain scarce and expensive to acquire (2023; 2025), leading to public datasets that are limited in size and suffer from highly imbalanced distributions (2022; 2024; 2025), which severely restricts model generalization. In addition, a substantial domain gap exists between experimental and computational spectra due to complex measurement conditions (2022), hindering the deployment of models trained on theoretical data. Furthermore, spectroscopy is inherently multimodal: it encompasses various spectral types (e.g., infrared, Raman, nuclear magnetic resonance) represented as either 1D signals or 2D images, often requiring integration with other molecular modalities such as molecular graphs, SMILES strings, and 3D conformations (2021; 2024). The heterogeneous nature and semantics of these data modalities pose significant challenges for deep learning systems. Finally, the field lacks standardized benchmarks, with a fragmented landscape of tasks and datasets making it difficult to systematically evaluate and compare model performance.
@@
To address these challenges, we introduce SpectrumLab, a modular platform that streamlines the entire lifecycle of AI-driven spectroscopy from data preprocessing to model evaluation. Built atop SpectrumLab, we construct SpectrumBench, a unified benchmark suite designed to evaluate machine learning models across diverse spectroscopic tasks and modalities. In contrast to existing approaches such as DiffSpectra (2025b) and MolSpectra (2025a), which rely on contrastive learning and diffusion architectures, we are among the first to incorporate multi-modal large language models (MLLMs) into spectroscopic learning, using their alignment capabilities to bridge heterogeneous data modalities.
Contributions.Our main contributions are:
(1) We present SpectrumLab, a standardized and extensible framework that integrates core modules for spectrum data processing, development for multimodal large language models, automatic result assessment, and visualization.
(2) We propose SpectrumAnnotator, a novel component that automatically generates task-specific benchmarks from seed datasets, enabling rapid prototyping and model testing.
(3) We develop SpectrumBench, a comprehensive benchmark suite spanning various spectroscopic tasks and modalities, equipped with unified evaluation protocols and public leaderboards to promote reproducibility and fair comparison.
@@ 2 Related Work
Figure 1: Representative SpectraML methods categorized by Spectral Type (left Y-axis) and Model Type (right Y-axis). Each dot indicates the use of a specific spectral modality or model architecture in a given method. Note that Raman is not included; thus, methods using it (e.g., DeepCID (2019)) are not shown on the left Y-axis.
@@
Machine Learning for Spectroscopy. Spectroscopy is fundamental for molecular structure analysis and scientific discovery, enabling insights into chemical properties and interactions (2025). Its applications span diverse scientific domains, including chemistry, material science, and drug development (2025; 2025). Machine learning techniques have been extensively applied in spectroscopy for tasks such as molecular structure elucidation (spectrum-to-molecule) (2008; 2018; 2018; 2019; 2019; 2020; 2020; 2021; 2024; 2025; 2025) and spectral simulation (molecule-to-spectral) (2017b; 2017; 2017; 2021; 2021; 2021; 2024). As illustrated in Figure 1, recent efforts have explored a variety of spectral modalities, such as IR (2024), NMR (2024), UV-Vis (2018), MS (2021), and Raman (2019), and have adopted heterogeneous deep learning model architectures, ranging from MLPs (2024) and CNNs (2024) to GNNs (2021) and Transformers (2024). Despite these rapid progresses, existing methods still face several limitations: (1) most studies are constrained to a single modality (e.g., IR or MS), lacking generalization across spectral types (2024); (2) the field lacks unified benchmarks and evaluation protocols, making objective comparisons difficult; (3) dataset sizes remain limited and imbalanced, further impeding reproducibility and robustness; (4) previous benchmarks does not support multi-modal large language models. These limitations highlight the need for standardized, cross-modal frameworks to advance machine learning for spectroscopy, especially spectroscopy foundation models.
Spectroscopy Foundation Models. While foundation models have shown promising progress in scientific discovery (2025; 2025), spectroscopy foundation models are still underexplored. This is largely due to the inherent multimodal nature of spectroscopic data, which combines spectral signals with diverse molecular representations. Although recent efforts such as SpectraFM (2024) and LSM1-MS2 (2024) have introduced pre-trained foundation models on Stellar and MS spectra for chemical property prediction, these models remain fundamentally single-modal, focusing solely on spectral information. Despite these challenges, the integration of spectroscopy into the foundation model paradigm holds significant promise for advancing automated analysis and multi-modal scientific discovery in the future.
@@
GR PA FM MR NovoBench (2024a) MolPuzzle (2024b) Multimodal Spec (2024) MassSpecGym (2024a) NMRNet (2025) ViBench (2025) SpectrumBench Ours
@@
Benchmark and Toolkits for Spectroscopy. Several benchmarks and toolkits have been developed to support spectroscopic machine learning research (2023; 2024b; 2024b; 2024b; 2024; 2024; 2025). However, many of these efforts remain limited in scope (either spectrum modalities or tasks), lacking extensibility and comprehensive evaluation across diverse spectroscopic tasks and modalities. For example, MassSpecGym (2024b) focuses solely on MS data and does not incorporate language descriptions, hindering support for multi-modal inputs. Although MolPuzzle (2024b) enables multi-modal inputs, it omits Raman spectra and lacks support for pure spectral understanding tasks. Furthermore, several toolkits (2024b; 2024b) do not provide interfaces for multi-modal large language models (MLLMs), and even MolPuzzle lacks benchmarking for more recent MLLMs. In contrast, our SpectrumLab is a unified, extensible, and reproducible platform that addresses these limitations by supporting a wide range of spectroscopic tasks, modalities, and integration with MLLMs. Table 1 systematically compares representative studies in terms of their spectral modality and task coverage. SpectrumLab not only fills critical gaps in data, evaluation, and tooling, but also establishes a new standard for spectroscopic AI and enables future advances in multi-modal, large-model-driven scientific discovery.
@@ 3 SpectrumBench
Overview
SpectrumBench is a unified benchmark suite for deep learning in spectroscopy, covering four hierarchical levels and 14 sub-tasks that span from spectroscopy understanding to generation. All questions and tasks are initially defined by domain experts, and subsequently refined and validated through expert review and rigorous quality assurance processes to ensure correctness and high quality. Compared to existing benchmarks, SpectrumBench offers broad modality and task coverage within a standardized, extensible framework for fair and reproducible model evaluation.
@@
Category Task Abbr. # questions Signal Spectrum Type Classification TC 55 Spectrum Quality Assessment QE 60 Basic Feature Extraction FE 51 Impurity Peak Detection ID 28 Perception Functional Group Recognition GR 45 Elemental Compositional Prediction EP 36 Peak Assignment PA 38 Basic Property Prediction PP 34 Semantic Molecular Structure Elucidation SE 80 Fusing Spectroscopic Modalities FM 39 Multimodal Molecular Reasoning MR 37 Generation Forward Problems FP 30 Inverse Problems IP 20 De Novo Generation DnG 19
Spectroscopic Type. Unlike previous benchmarks that are limited to a single spectroscopic modality or narrowly defined data types (2024a), SpectrumBench integrates a diverse array of spectroscopic data sources. Our SpectrumBench benchmark currently includes more than 10 distinct types of spectroscopic data, such as infrared (IR), nuclear magnetic resonance (NMR), and mass spectrometry (MS). As illustrated in Figure 7(see Appendix A), this comprehensive data foundation accurately reflects the diverse and complex multi-modal spectroscopic scenarios encountered in real-world applications.
Task. In contrast to previous benchmarks that primarily focus on molecule elucidation or spectrum simulation, SpectrumBench encompasses a much broader spectrum of task types. SpectrumBench is organized according to a multi-level hierarchical taxonomy that systematically covers tasks ranging from low-level signal analysis to high-level semantic reasoning and generative challenges. This taxonomy, developed through expert consultation and iterative refinement, comprises four principal layers: signal, perception, semantic, and generation. Each layer is further divided into several subcategories, capturing a diverse set of scientific and application-driven tasks. Detailed definitions and representative examples for each task layer are provided in the Appendix A.
Data Curation Pipeline
@@
Task Construction.Spectroscopic machine learning encompasses a wide spectrum of tasks, driven by the intrinsic complexity of molecular structures and the multifaceted nature of spectroscopic data. These tasks often involve diverse input modalities (e.g., molecular graphs, SMILES, textual prompts) and equally varied outputs (e.g., spectra, chemical attributes, structured predictions), which reflect the real-world demands of chemical analysis, property reasoning, and molecular generation. To illustrate this diversity, we organize existing spectroscopic tasks into four broad input-output categories:
(1) Molecule-to-Spectrum(Spectrum Simulation) aims at generating a spectrum based on molecular structure.
(2) Spectrum-to-Molecule (Molecule Elucidation) refers to the tasks that infer molecular structures from spectra.
(3) Text-to-Any111“Any” encompasses various data modalities, such as molecular representations, spectral data, or peptide sequences.(De novo Generation) refers to the task of generating novel, diverse, and reasonable molecular structures (SMILES string, 2D molecular graph) and/or predicting multimodal information (spectra, properties) according to specific goals (e.g., molecules of a specific nature, ligands of a specific target).
Moreover, in previous studies, many tasks involving inferring molecular structures from spectra were also categorized under “de novo generation”(2024b; 2025). While this has some rationality, for the sake of consistency in our task framework, we clarify that our defined de novo generation task has distinct characteristics: its input consists solely of textual descriptions, which may include specifications of molecular properties(e.g., desired chemical natures, target-binding affinities), without involving spectral data as input. Meanwhile, the output scope is broader, encompassing not only molecular structures but also spectral and textual descriptions of molecules.
(4) Any-to-Text (Understanding). Tasks in signal, perception, semantic layers fall under the “Understanding” category. Its task type is presented in the form of a multiple-choice questions, which may include tasks such as inferring the molecular structure from spectrum (e.g., functional group recognition, peak attribution tasks). This partially overlaps with the molecular elucidation tasks described above. For a compromise design, we use the output form to distinguish between them. The question format of “Understanding” tasks will only be multiple-choice questions, which means the output is text.
@@
Taxonomy Definition. These input-output patterns offer a high-level overview of the task landscape. However, prior works often cover only a subset of them, limiting both their generalizability and their ability to benchmark diverse ML capabilities. We show these patterns in Table 1, which highlights substantial heterogeneity across existing methods. To address this limitation and support more structured, extensible benchmarking, we propose a four-level hierarchical taxonomy tailored to spectroscopic machine learning: Signal, Perception, Semantic, and Generation—is designed to reflect the logic of real-world scientific workflows in spectroscopy. As depicted in Figure 2, this layered structure systematically provides a robust framework for our 14 meticulously designed tasks detailed in Table 2. (1) Signal level: This foundational layer focuses on the direct analysis and processing of raw spectral data, such as spectrum type classification and peak detection. Tasks at this level are designed to extract and refine primary features from experimental measurements, mirroring the initial steps taken by chemists to prepare and interpret spectra in the laboratory. This level primarily encompasses Any-to-Text(Understanding) tasks that operate directly on raw signal data.
@@
Seed Data Preparation. The seed datasets used in this work are curated from three primary sources to ensure both diversity and scientific rigor. (1) Proprietary collections and in-house experimental data: These include unpublished spectroscopic measurements and curated datasets generated within our collaborating laboratories, providing access to high-quality, domain-specific spectra. This source comprises approximately 238,869 molecular data points covering 8 types of spectra. Being experimental spectra, these data offer higher authenticity and usability compared to most computationally generated spectra. (2) Public repositories and benchmark datasets: We draw upon data from a range of widely recognized and authoritative sources, including SDBS (2025), QM9S (2023), NovoBench (2024a), and MolPuzzle (2024b), among others. In total, seven distinct repositories and public datasets are integrated, collectively encompassing over 1.01 million unique chemical compounds. This comprehensive aggregation ensures both the diversity and representativeness of real-world spectroscopic scenarios. (3) Literature mining: To further expand the dataset, we systematically extract spectral data from the Supporting Information sections of peer-reviewed publications, specifically focusing on articles published in leading journals such as Journal of the American Chemical Society (JACS) and ACS Catalysis. This multi-source curation strategy guarantees that the seed datasets are both diverse and representative, laying a robust foundation for subsequent benchmark construction and model evaluation.
@@
As illustrated in Figure 2, the construction of seed datasets begins with the aggregation of raw data from multiple authoritative repositories. All collected datasets undergo a unified data processing pipeline, which systematically maps each entry into three core chemical spaces: SMILES string, molecular formula, and spectra. Through rigorous cleaning, normalization, and deduplication, we ensure the consistency and reliability of the data. Human annotation and alignment are then performed to guarantee scientific accuracy and completeness. The resulting seed datasets are organized at the level of individual chemical substances, with each record containing the compound’s SMILES, molecular formula, and a structured set of associated spectra, all stored in a standardized JSON format to facilitate downstream annotation and interoperability. Detailed descriptions of the seed datasets and the standardization process are provided in Appendix F.
@@
Data Annotation. We use two annotation methods: automated and manual annotation. (1) Automated annotation (SpectrumAnnotator). For tasks characterized by well-defined rules and moderate complexity—such as spectrum recognition, basic feature extraction from spectrum, and other standard spectroscopic benchmarks—we design SpectrumAnnotator, a core contribution of this work. SpectrumAnnotatoris a novel, self-developed annotation framework that harnesses the zero-shot and multi-modal reasoning capabilities of state-of-the-art MLLMs. Given curated seed datasets and a set of pre-defined benchmark prompts, SpectrumAnnotator  automatically designs and generates high-quality, multi-modal benchmark data, including both image-text pairs and complex reasoning tasks. This automated pipeline enables rapid, scalable, and consistent construction of diverse benchmarks, crucially streamlining annotation and facilitating the creation of challenging multi-modal tasks essential for next-generation model development. Further technical details and implementation specifics of SpectrumAnnotator are provided in Appendix B. (2) Manual Annotation. For more complex or open-ended tasks, particularly those involving multi-step reasoning or sophisticated scientific interpretation, manual annotation by domain experts is indispensable. Human annotators ensure the scientific validity and depth of the benchmark, especially in cases where automated methods cannot handle. Throughout the annotation process, we emphasize human-AI collaboration to maximize both efficiency and accuracy.

### v4

@@ Abstract
Deep learning holds immense promise for spectroscopy, yet research and evaluation in this emerging field often lack standardized formulations. To address this issue, we introduce SpectrumLab, a pioneering unified platform designed to systematize and accelerate deep learning research in spectroscopy. SpectrumLab integrates three core components: a comprehensive Python library featuring essential data processing and evaluation tools, along with leaderboards; an innovative SpectrumAnnotator module that generates high-quality benchmarks from limited seed data; and SpectrumBench, a multi-layered benchmark suite covering 14 spectroscopic tasks and over 10 spectrum types, featuring spectra curated from over 1.2 million distinct chemical substances. Thorough empirical studies on SpectrumBench with 23 cutting-edge multimodal LLMs reveal critical limitations of current approaches. We hope SpectrumLab will serve as a crucial foundation for future advancements in deep learning-driven spectroscopy. The anonymous code are available at https://github.com/little1d/SpectrumLab.
@@ 1 Introduction
Spectroscopy, which investigates the interaction between electromagnetic radiation and matter, provides a powerful way to investigate the molecular structure and properties (Elias et al., 2004; Prasad et al., 2025). By capturing characteristic patterns, such as peaks and shifts, in signals analogous to audio waveforms, spectroscopy offers a compact, information-rich representation of molecular systems (Ralbovsky & Lednev, 2020). This low-dimensional encoding is indispensable in chemistry (Silber et al., 2016; Seo et al., 2017), and life sciences (Ralbovsky & Lednev, 2020; Zhang et al., 2023; Gasparin et al., 2025). It is not only central to molecular structure elucidation (i.e., Spectrum-to-Molecule structure) and property prediction, but also a key enabler for new material discovery and drug screening. In recent years, machine learning methods, especially deep learning, have demonstrated tremendous potential in spectroscopic data analysis, opening a new era of automation and intelligence in spectroscopy research (Gastegger et al., 2017b; Gerrard et al., 2019; Fine et al., 2020; Han et al., 2022; Zou et al., 2023; Devata et al., 2024; Lu et al., 2025).
@@
Despite recent advances, deep learning for spectroscopy still faces several fundamental challenges. Specifically, high-quality experimental spectral data remain scarce and expensive to acquire (van de Sande et al., 2023; Flanagan et al., 2025), leading to public datasets that are limited in size and suffer from highly imbalanced distributions (Bongiorno et al., 2022; Stenning et al., 2024; Peng et al., 2025), which severely restricts model generalization. In addition, a substantial domain gap exists between experimental and computational spectra due to complex measurement conditions (Agarwala et al., 2022), hindering the deployment of models trained on theoretical data. Furthermore, spectroscopy is inherently multimodal: it encompasses various spectral types (e.g., infrared, Raman, nuclear magnetic resonance) represented as either 1D signals or 2D images, often requiring integration with other molecular modalities such as molecular graphs, SMILES strings, and 3D conformations (Litsa et al., 2021; Devata et al., 2024). The heterogeneous nature and semantics of these data modalities pose significant challenges for deep learning systems. Finally, the field lacks standardized benchmarks, with a fragmented landscape of tasks and datasets making it difficult to systematically evaluate and compare model performance.
@@
To address these challenges, we introduce SpectrumLab, a modular platform that streamlines the entire lifecycle of AI-driven spectroscopy from data preprocessing to model evaluation. Built on top of SpectrumLab, we construct SpectrumBench, a unified benchmark suite designed to evaluate machine learning models across diverse spectroscopic tasks and modalities. In contrast to existing approaches such as DiffSpectra (Wang et al., 2025b) and MolSpectra (Wang et al., 2025a), which rely on contrastive learning and diffusion architectures, we are among the first to incorporate multi-modal large language models (MLLMs) into spectroscopic learning, using their alignment capabilities to bridge heterogeneous data modalities.
Figure 1: Overview of SpectrumBench.
Our main contributions are: (1) We introduce SpectrumLab, the first standardized framework tailored for spectroscopic machine learning with multimodal large language models, enabling reproducible pipelines from raw spectra to evaluation. (2) We design SpectrumAnnotator, an automatic benchmark generator that constructs task-specific datasets from spectrum seeds, greatly accelerating prototyping and stress-testing of new models. (3) We release SpectrumBench, a large-scale benchmark suite covering diverse spectroscopic modalities and tasks, accompanied by unified evaluation protocols and public leaderboards to foster fair comparison and community progress.
@@ 2 Related Work
Figure 2: Representative SpectraML methods categorized by Spectral Type (left Y-axis) and Model Type (right Y-axis). Each dot indicates the use of a specific spectral modality or model architecture in a given method. Note that Raman is not included; thus, methods using it (e.g., DeepCID (Fan et al., 2019)) are not shown on the left Y-axis.
@@
Machine Learning for Spectroscopy. Spectroscopy is fundamental for molecular structure analysis and scientific discovery, enabling insights into chemical properties and interactions (Guo et al., 2025). Its applications span diverse scientific domains, including chemistry, material science, and drug development (Shao et al., 2025; Sun et al., 2025). Machine learning techniques have been extensively applied in spectroscopy for tasks such as molecular structure elucidation (spectrum-to-molecule) (Kuhn et al., 2008; De Vijlder et al., 2018; Paruzzo et al., 2018; Fan et al., 2019; Nguyen et al., 2019; Ji et al., 2020; Fine et al., 2020; Huang et al., 2021; Hu et al., 2024; Lu et al., 2025; Liu et al., 2025) and spectral simulation (molecule-to-spectral) (Gastegger et al., 2017a; Zhou et al., 2017; Liu et al., 2017; McGill et al., 2021; Guan et al., 2021; Ren et al., 2021; Young et al., 2024). As illustrated in Figure 2, recent efforts have explored a variety of spectral modalities, such as IR (Hu et al., 2024), NMR (Hu et al., 2024), UV-Vis (De Vijlder et al., 2018), MS (Huber et al., 2021), and Raman (Fan et al., 2019), and have adopted heterogeneous deep learning model architectures, ranging from MLPs (Stienstra et al., 2024) and CNNs (Alberts et al., 2024b) to GNNs (McGill et al., 2021) and Transformers (Alberts et al., 2024a). Despite these rapid progresses, existing methods still face several limitations: (1) most studies are constrained to a single modality (e.g., IR or MS), lacking generalization across spectral types (Beck et al., 2024); (2) the field lacks unified benchmarks and evaluation protocols, making objective comparisons difficult; (3) dataset sizes remain limited and imbalanced, further impeding reproducibility and robustness; (4) previous benchmarks does not support multi-modal large language models. These limitations highlight the need for standardized, cross-modal frameworks to advance machine learning for spectroscopy, especially spectroscopy foundation models.
Spectroscopy Foundation Models. While foundation models have shown promising progress in scientific discovery (Tan et al., 2025; Xia et al., 2025), spectroscopy foundation models are still underexplored. This is largely due to the inherent multimodal nature of spectroscopic data, which combines spectral signals with diverse molecular representations. Although recent efforts such as SpectraFM (Koblischke & Bovy, 2024) and LSM1-MS2 (Asher et al., 2024) have introduced pre-trained foundation models on Stellar and MS spectra for chemical property prediction, these models remain fundamentally single-modal, focusing solely on spectral information. Despite these challenges, the integration of spectroscopy into the foundation model paradigm holds significant promise for advancing automated analysis and multi-modal scientific discovery in the future.
@@
GR PA FM MR NovoBench (Zhou et al., 2024a) MolPuzzle (Guo et al., 2024b) Multimodal Spec (Alberts et al., 2024b) MassSpecGym (Bushuiev et al., 2024b) NMRNet (Xu et al., 2025) ViBench (Lu et al., 2025) SpectrumBench Ours
@@
Benchmark and Toolkits for Spectroscopy. Several benchmarks and toolkits have been developed to support spectroscopic machine learning research (Heid et al., 2023; Zhou et al., 2024b; Bushuiev et al., 2024a; Guo et al., 2024b; Devata et al., 2024; Ruan et al., 2024; Guo et al., 2025). However, many of these efforts remain limited in scope (either spectrum modalities or tasks), lacking extensibility and comprehensive evaluation across diverse spectroscopic tasks and modalities. For example, MassSpecGym (Bushuiev et al., 2024a) focuses solely on MS data and does not incorporate language descriptions, hindering support for multi-modal inputs. Although MolPuzzle (Guo et al., 2024b) enables multi-modal inputs, it omits Raman spectra and lacks support for pure spectral understanding tasks. Furthermore, several toolkits (Bushuiev et al., 2024a; Zhou et al., 2024b) do not provide interfaces for multi-modal large language models (MLLMs), and even MolPuzzle lacks benchmarking for more recent MLLMs. In contrast, our SpectrumLab is a unified, extensible, and reproducible platform that addresses these limitations by supporting a wide range of spectroscopic tasks, modalities, and integration with MLLMs. Table 1 systematically compares representative studies in terms of their spectral modality and task coverage. SpectrumLab not only fills critical gaps in data, evaluation, and tooling, but also establishes a new standard for spectroscopic AI and enables future advances in multi-modal, large-model-driven scientific discovery.
@@ 3 SpectrumBench
Overview. SpectrumBench is a unified benchmark suite for deep learning in spectroscopy, covering four hierarchical levels and 14 sub-tasks that span from spectroscopy understanding to generation. All questions and tasks are initially defined by domain experts, and subsequently refined and validated through expert review and rigorous quality assurance processes. Compared to existing benchmarks, SpectrumBench offers broad modality and task coverage within a standardized, extensible framework for fair and reproducible model evaluation.
@@
Category Task # questions Signal Spectrum Type Classification (TC) 55 Spectrum Quality Assessment (QE) 60 Basic Feature Extraction (FE) 51 Impurity Peak Detection (ID) 28 Perception Functional Group Recognition 45 Elemental Compositional Prediction (EP) 36 Peak Assignment (PA) 38 Basic Property Prediction (PP) 34 Semantic Molecular Structure Elucidation (SE) 80 Fusing Spectroscopic Modalities (FM) 39 Multimodal Molecular Reasoning (MR) 37 Generation Forward Problems (FP) 30 Inverse Problems (IP) 20 De Novo Generation (DnG) 19
Spectroscopic Type. Unlike previous benchmarks that are limited to a single spectroscopic modality or narrowly defined data types (Bushuiev et al., 2024b), SpectrumBench integrates a diverse array of spectroscopic data sources. Our SpectrumBench benchmark currently includes more than 10 distinct types of spectroscopic data, such as infrared (IR), nuclear magnetic resonance (NMR), and mass spectrometry (MS). As illustrated in Figure 1, this comprehensive data foundation accurately reflects the diverse and complex multi-modal spectroscopic scenarios encountered in real-world applications.
Task. In contrast to previous benchmarks that primarily focus on molecule elucidation or spectrum simulation, SpectrumBench encompasses a much broader spectrum of task types. SpectrumBench is organized according to a multi-level hierarchical taxonomy that systematically covers tasks ranging from low-level signal analysis to high-level semantic reasoning and generative challenges. This taxonomy, developed through expert consultation and iterative refinement, comprises four principal layers: signal, perception, semantic, and generation. Each layer is further divided into several subcategories, capturing a diverse set of scientific and application-driven tasks. Detailed definitions and representative examples for each task layer are provided in the Appendix B.
3.1 Data Curation Pipeline
@@
Task Construction. Spectroscopic machine learning encompasses a wide spectrum of tasks, driven by the intrinsic complexity of molecular structures and the multifaceted nature of spectroscopic data. These tasks often involve diverse input modalities (e.g., molecular graphs, SMILES, textual prompts) and equally varied outputs (e.g., spectra, chemical attributes, structured predictions), which reflect the real-world demands of chemical analysis, property reasoning, and molecular generation. To illustrate this diversity, we organize existing spectroscopic tasks into four broad input-output categories: (1) Molecule-to-Spectrum (Spectrum Simulation) aims at generating a spectrum based on molecular structure. (2) Spectrum-to-Molecule (Molecule Elucidation) refers to the tasks that infer molecular structures from spectra. (3) Text-to-Any111“Any” encompasses various data modalities, such as molecular representations, spectral data, or peptide sequences. (De novo Generation) refers to the task of generating novel, diverse, and reasonable molecular structures (SMILES string, 2D molecular graph) and/or predicting multimodal information (spectra, properties) according to specific goals (e.g., molecules of a specific nature, ligands of a specific target). Moreover, in previous studies, many tasks involving inferring molecular structures from spectra were also categorized under “de novo generation” (Bushuiev et al., 2024a; Lu et al., 2025). While this has some rationality, for the sake of consistency in our task framework, we clarify that our defined de novo generation task has distinct characteristics: its input consists solely of textual descriptions, which may include specifications of molecular properties (e.g., desired chemical natures, target-binding affinities), without involving spectral data as input. Meanwhile, the output scope is broader, encompassing not only molecular structures but also spectral and textual descriptions of molecules. (4) Any-to-Text (Understanding). Tasks in signal, perception, semantic layers fall under the “Understanding” category. Its task type is presented in the form of multiple-choice questions, which may include tasks such as inferring the molecular structure from spectrum (e.g., functional group recognition, peak attribution tasks). This partially overlaps with the molecular elucidation tasks described above. For a compromise design, we use the output form to distinguish between them. The question format of “Understanding” tasks will only be multiple-choice questions, which means the output is text.
@@
Taxonomy Definition. These input-output patterns offer a high-level overview of the task landscape. However, previous works often cover only a subset of them, limiting both their generalizability and their ability to benchmark diverse ML capabilities. We show these patterns in Table 1, which highlights substantial heterogeneity between existing methods. To address this limitation and support more structured and extensible benchmarking, we propose a four-level hierarchical taxonomy tailored to spectroscopic machine learning: Signal, Perception, Semantic, and Generation—is designed to reflect the logic of real-world scientific workflows in spectroscopy. As depicted in Figure 3, this layered structure systematically provides a robust framework for our 14 meticulously designed tasks detailed in Table 2.
(1) Signal level: This foundational layer focuses on the direct analysis and processing of raw spectral data, such as spectrum type classification and peak detection. Tasks at this level are designed to extract and refine primary features from experimental measurements, mirroring the initial steps taken by chemists to prepare and interpret spectra in the laboratory. This level primarily encompasses Any-to-Text(Understanding) tasks that operate directly on raw signal data.
@@
Figure 4: SpectrumLab is the main component of SpectrumWorld, which could be extended conveniently. More multi-modal spectrum data could be included, as well as MLLMs.
@@
Seed Data Preparation. The seed datasets for this work were curated from three primary sources: proprietary experimental data, public repositories, and literature mining. As illustrated in Figure 3, the construction of seed datasets begins with the aggregation of raw data from multiple sources. All collected datasets undergo a unified processing pipeline that systematically maps each entry into three core chemical spaces: SMILES string, molecular formula, and spectra. Rigorous cleaning, normalization, and deduplication ensure data consistency and reliability. Following this, human annotation and alignment are performed to guarantee scientific accuracy and completeness. The resulting seed datasets are organized at the level of individual chemical substances, with each record containing the compound’s SMILES, molecular formula, and a structured set of associated spectra, all stored in a standardized JSON format to facilitate downstream annotation. Detailed descriptions of the seed datasets and the standardization process are provided in Appendix G. We ensure that the curated seed data are not contaminated.
@@
Data Annotation. We use two annotation methods: automated and manual annotation. (1) Automated annotation (SpectrumAnnotator). For tasks characterized by well-defined rules like spectrum recognition , we design SpectrumAnnotator, a core contribution of this work. SpectrumAnnotatoris a novel, self-developed annotation framework that harnesses the zero-shot and multi-modal reasoning capabilities of state-of-the-art MLLMs. Given curated seed datasets and a set of pre-defined benchmark prompts, SpectrumAnnotator automatically designs and generates high-quality, multi-modal benchmark data. Further technical details and implementation specifics of SpectrumAnnotator are provided in Appendix C. (2) Manual Annotation. For more complex or open-ended tasks, particularly those involving multi-step reasoning or sophisticated scientific interpretation, manual annotation by domain experts is indispensable. Human annotators ensure the scientific validity and depth of the benchmark, especially in cases where automated methods cannot handle.